# Irreducibility of three dimensional automorphic Galois representations over totally real fields

### v1

@@ 1. Introduction
Let $K$ be a totally real (or CM) number field, $\pi$ a regular algebraic cuspidal automorphic representation of $\mathrm{GL}_{n}(\mathbb{A}_{K})$, and $\iota:\mathbb{C}\stackrel{{\scriptstyle\sim}}{{\rightarrow}}\overline{\mathbb{Q}}_{\ell}$ a field isomorphism. Denote by $\operatorname{Gal}_{K}$ the absolute Galois group of $K$, by $\Sigma_{K}$ the set of finite places of $K$, by $S_{\ell}\subset\Sigma_{K}$ the subset of places $v$ that divide $\ell$, and by $S\subset\Sigma_{K}$ the subset of places $v$ such that the local component $\pi_{v}$ is ramified. Attached to the pair $(\pi,\iota)$ is a semisimple (conjecturally irreducible, see e.g., [Ram08]) $\ell$-adic representation of $K$
@@
When $K$ is totally real and $n=3$ (resp. $4$), D. Ramakrishnan obtained (by using an L-function argument) the irreducibility of (1) assuming some $\ell$-adic Hodge properties [Ram13, Theorem B]. In this article, we prove such irreducibility unconditionally (Theorem 1.1) and deduce in return, the properties (dR) and (Cr) above, if $\rho_{\ell}$ is of irreducible type ($A_{2}$) and $\ell$ belongs to a Dirichlet density one set of rational primes (Theorem 3.6).
@@
The new input for Theorem 1.1 is the following result which is of independent interest.
Theorem 1.2.
Let $K$ and $E\subset\overline{\mathbb{Q}}_{\ell}$ be number fields, and $\rho_{\ell}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{\mathbb{Q}}_{\ell})$ a semisimple $E$-rational $\ell$-adic representation of $K$. If $\psi_{\ell}$ is a weak abelian direct summand of $\rho_{\ell}$, then it is locally algebraic, and in particular de Rham.
We briefly describe the remaining sections of this article. We first recall basic notions and facts about (abelian) $\ell$-adic representations in §$\mathsection 2.1$§$\mathsection 2.5$. Then we introduce weak abelian direct summands (a notion weaker than abelian subrepresentations) of an almost everywhere unramified semisimple $\ell$-adic Galois representation and completely characterize them in case the algebraic monodromy is connected (Proposition 2.9) in §$\mathsection 2.6$. Theorem 1.2 is established in §$\mathsection 2.7$ by using §$\mathsection 2.1$§$\mathsection 2.6$ and a remarkable result of Waldschmidt. Finally in §$\mathsection 2.8$§$\mathsection 2.9$, we formulate a conjecture on $E$-rationality of the abelian part $\rho_{\ell}^{\text{ab}}$ (Conjecture 2.14) with some evidence (Theorem 2.15) and prove the compatibility of the weak abelian part of a connected semisimple Serre compatible system (Theorem 2.20).
@@ 2.3. Local algebraicity
Moreover, $\phi_{\ell}$ is locally algebraic if and only if these local representations are Hodge-Tate, see [Se98, Appendix of Chap. III].
@@ 2.4. Equivalence of three notions
The direction ($E$-SCS) $\Rightarrow$ ($E$-rat) is obvious. Finally for ($E$-rat) $\Rightarrow$ (Loc-alg), according to the strategy of Serre in [Se98, Chap. III.3] there are two steps: the first one is to show that $\phi_{\ell}$ is almost locally algebraic (i.e., $\phi_{\ell}^{N}$ is locally algebraic for some $N\in\mathbb{N}$) and the second one is to show that almost local algebraicity implies local algebraicity for an $E$-rational $\phi_{\ell}$ (see Proposition 2.12). The first step is a direct consequence of Theorem 2.2 below which is proven (see e.g., [He82, Theorem 7]) by a remarkable theorem of Waldschmidt on transcendental numbers [Wa81].
@@ 2.9. Abelian and weak abelian part of compatible system
Let $\{\rho_{\lambda}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{E}_{\lambda})\}$ be a semisimple Serre compatible system of $K$ defined over $E$ unramified outside a finite $S\subset\Sigma_{K}$. Let $\mathbf{G}_{\lambda}$ be the algebraic monodromy group of $\rho_{\lambda}$. The connectedness and formal character of $\mathbf{G}_{\lambda}$ is independent of $\lambda$; see [Se81b]. We say that $\{\rho_{\lambda}\}$ is connected if $\mathbf{G}_{\lambda}$ is connected for some $\lambda$ (hence all $\lambda$). Define $\{\rho_{\lambda}^{\text{ab}}\}$ (resp. $\{\rho_{\lambda}^{\text{wab}}\}$, well-defined if $\{\rho_{\lambda}\}$ is connected by Proposition 2.9) as the abelian part (resp. weak abelian part) of the Serre compatible system $\{\rho_{\lambda}\}$. One can ask if $\{\rho_{\ell}^{\text{ab}}\}$ (resp. $\{\rho_{\lambda}^{\text{wab}}\}$) is still a Serre compatible system defined over $E$ when it is non-zero. This question is studied in [Hu20] (in case $E=\mathbb{Q}$) for the abelian part by using some $\lambda$-independence results of $\mathbf{G}_{\lambda}$, such as the following.
@@ 3.2. Irreducibility and algebraic monodromy groups
Suppose $\det(\sigma_{\lambda})\cdot\tau_{\lambda}^{-2}$ is trivial. Then $\rho_{\lambda}\otimes\tau_{\lambda}^{-1}$ is self-dual. Since $\tau_{\lambda}$ is locally algebraic, $\pi$ is essentially self-dual by multiplicity one [JS81]. But in this case, $\rho_{\lambda}$ is irreducible (see [Hu23a, §$\mathsection 4.4.3$]). We conclude that $\rho_{\lambda}$ is irreducible.
@@ Acknowledgments
G. Böckle acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Collaborative Research Centre TRR 326 Geometry and Arithmetic of Uniformized Structures, project number 444845124. C.-Y. Hui was partially supported by NSFC (no. 12222120) and a Humboldt Research Fellowship.

### v2

# Weak abelian direct summands and irreducibility of Galois representations

@@ 1. Introduction
1.1. Weak abelian direct summands
Let $K$ be a number field, $\overline{K}$ an algebraic closure of $K$, and $\operatorname{Gal}_{K}:=\operatorname{Gal}(\overline{K}/K)$ the absolute Galois group of $K$. Denote by $\Sigma_{K}$ the set of finite places of $K$ and by $S_{\ell}$ the subset of places that divide $\ell$. An $\ell$-adic representation of $K$
is said to be $E$-rational for some number field $E\subset\overline{\mathbb{Q}}_{\ell}$ if for almost all $v\in\Sigma_{K}$, the Galois representation $\rho_{\ell}$ is unramified at $v$ and the characteristic polynomial of $\rho_{\ell}(\mathrm{Frob}_{v})$ has coefficients in $E$. Such $E$-rational $\ell$-adic representations provide Galois theoretic incarnations of many objects in number theory and arithmetic geometry. For instance, Faltings isogeny theorem [Fa83] asserts that two abelian varieties defined over $K$ are isogenous if their attached ($\mathbb{Q}$-rational) $\ell$-adic representations of $K$ are equivalent.
[…]
Let $\psi_{\ell}$ be an abelian semisimple $\ell$-adic representation of $K$. We say that $\psi_{\ell}$ is a weak abelian direct summand of $\rho_{\ell}$ if there is a subset $\mathcal{L}\subset\Sigma_{K}$ of Dirichlet density one such that for all $v\in\mathcal{L}$, $\psi_{\ell}$ and $\rho_{\ell}$ are unramified at $v$ and the characteristic polynomial of $\psi_{\ell}(\mathrm{Frob}_{v})$ divides those of $\rho_{\ell}(\mathrm{Frob}_{v})$. Abelian direct summands (i.e., subrepresentations) of $\rho_{\ell}$ are obvious examples of weak abelian direct summands; if $n$ is odd and $\rho_{\ell}$ factors through $\mathrm{GO}_{n}=\mathrm{SO}_{n}\times\mathbb{G}_{m}$, then the homothety character of $\rho_{\ell}$ is also a weak abelian direct summand. The goals of this article are to investigate this new notion (more flexible than abelian subrepresentations) and give some applications to compatible system and automorphic Galois representations.
Under mild conditions on the algebraic monodromy group $\mathbf{G}_{\ell}$ of $\rho_{\ell}$ (e.g., when $\mathbf{G}_{\ell}$ is connected), we completely characterize the weak abelian direct summands of an $E$-rational semisimple $\rho_{\ell}$ (Proposition 2.9). This, together with Serre-Waldschmidt theory, lead to the main theorem below, which generalizes the implication
$E$-rationality   $\Rightarrow$  local algebraicity
for abelian semisimple $\rho_{\ell}$ to the general (i.e., non-abelian) situation.
Theorem 1.1.
Let $K$ and $E\subset\overline{\mathbb{Q}}_{\ell}$ be number fields, and $\rho_{\ell}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{\mathbb{Q}}_{\ell})$ a semisimple $E$-rational $\ell$-adic representation of $K$. If $\psi_{\ell}$ is a weak abelian direct summand of $\rho_{\ell}$, then it is locally algebraic, and thus its local representations at places above $\ell$ are de Rham.
Let $E$ be a number field and $\{\rho_{\lambda}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{E}_{\lambda}):~{}\lambda\in\Sigma_{E}\}$ a semisimple (Serre) compatible system of $K$ defined over $E$. Denote the abelian part (i.e., the maximal abelian subrepresentation) of $\rho_{\lambda}$ by $\rho_{\lambda}^{\text{ab}}$. It is conjectured that the family $\{\rho_{\lambda}^{\text{ab}}\}$ is a compatible system; but this seems a difficult question that is only resolved in some cases, cf. [Hu20] Suppose the algebraic monodromy group $\mathbf{G}_{\lambda}$ of $\rho_{\lambda}$ is connected for all $\lambda$. The weak abelian part of $\rho_{\lambda}$, denoted by $\rho_{\lambda}^{\text{wab}}$, is defined to be the maximal weak abelian direct summand of $\rho_{\lambda}$. As an application of our results on weak abelian direct summands, we show that $\{\rho_{\lambda}^{\text{wab}}\}$ is likewise a compatible system (Theorem 2.20). In next subsection, we describe our results on some three dimensional automorphic Galois representations.
1.2. Automorphic Galois representations
Let $K$ be a totally real (or CM) number field, $\pi$ a regular algebraic cuspidal automorphic representation of $\mathrm{GL}_{n}(\mathbb{A}_{K})$, and $\iota:\mathbb{C}\stackrel{{\scriptstyle\sim}}{{\rightarrow}}\overline{\mathbb{Q}}_{\ell}$ a field isomorphism. Denote by $S\subset\Sigma_{K}$ the subset of places $v$ such that the local component $\pi_{v}$ is ramified. Attached to the pair $(\pi,\iota)$ is a semisimple (conjecturally irreducible, see e.g., [Ram08]) $\ell$-adic representation of $K$
@@ 1.2. Automorphic Galois representations
When $K$ is totally real and $n=3,4$, D. Ramakrishnan obtained the irreducibility of (1) assuming some $\ell$-adic Hodge properties [Ram13, Theorem B]. The idea for $n=3$ is as follows. If on the contrary $\rho_{\pi,\iota}$ is reducible, then it contains a one-dimensional subrepresentation $\tau_{\lambda}$ (a weak abelian direct summand). Assuming $\tau_{\lambda}$ is Hodge-Tate, Ramakrishnan deduces a contradiction to the cuspidality of $\pi$ by an L-function argument. As a consequence of Theorem 1.1, we obtain the following unconditional irreducibility result, which inspires this article.
@@
In the opposite direction, we deduce in Theorem 3.7 for $\rho_{\ell}$ of irreducible type ($A_{2}$) that properties (dR) and (Cr) hold for rational primes $\ell$ in a Dirichlet density one set.
1.3. Structure of the article
We briefly describe the remaining sections of this article. We first recall basic notions and facts about (abelian) $\ell$-adic representations in §$\mathsection 2.1$§$\mathsection 2.5$. Then we introduce weak abelian direct summands of an almost everywhere unramified semisimple $\ell$-adic Galois representation and completely characterize them under mild conditions on $\mathbf{G}_{\ell}$ (Proposition 2.9) in §$\mathsection 2.6$. The main Theorem 1.1 is established in §$\mathsection 2.7$ by using §$\mathsection 2.1$§$\mathsection 2.6$ and Serre-Waldschmidt theory. Finally in §$\mathsection 2.8$§$\mathsection 2.9$, we formulate a conjecture on $E$-rationality of the abelian part $\rho_{\ell}^{\text{ab}}$ (Conjecture 2.14) with some evidence (Theorem 2.15) and prove the compatibility of the weak abelian part of a connected semisimple Serre compatible system (Theorem 2.20).
@@ 2.3. Local algebraicity
[…]
@@ 2.4. Equivalence of three notions
The direction ($E$-SCS) $\Rightarrow$ ($E$-rat) is obvious. Finally for ($E$-rat) $\Rightarrow$ (Loc-alg), according to the strategy of Serre in [Se98, Chap. III.3] there are two steps: the first one is to show that $\phi_{\ell}$ is almost locally algebraic (i.e., $\phi_{\ell}^{N}$ is locally algebraic for some $N\in\mathbb{N}$) and the second one is to show that almost local algebraicity implies local algebraicity for an $E$-rational $\phi_{\ell}$ (see Proposition 2.12). The first step is a direct consequence of Theorem 2.2 below which is due to Waldschmidt [Wa81] (see e.g., [He82, Theorem 7]).
@@ 2.9. Abelian and weak abelian part of compatible system
Let $\{\rho_{\lambda}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{E}_{\lambda})\}$ be a semisimple Serre compatible system of $K$ defined over $E$ unramified outside a finite $S\subset\Sigma_{K}$. Let $\mathbf{G}_{\lambda}$ be the algebraic monodromy group of $\rho_{\lambda}$. The connectedness and formal character of $\mathbf{G}_{\lambda}$ is independent of $\lambda$; see [Se81b]. We say that $\{\rho_{\lambda}\}$ is connected if $\mathbf{G}_{\lambda}$ is connected for some $\lambda$ (hence all $\lambda$). Define $\{\rho_{\lambda}^{\text{ab}}\}$ (resp. $\{\rho_{\lambda}^{\text{wab}}\}$, well-defined if $\{\rho_{\lambda}\}$ is connected by Proposition 2.9 and Corollary 2.10(i)) as the abelian part (resp. weak abelian part) of the Serre compatible system $\{\rho_{\lambda}\}$. One can ask if $\{\rho_{\ell}^{\text{ab}}\}$ (resp. $\{\rho_{\lambda}^{\text{wab}}\}$) is still a Serre compatible system defined over $E$ when it is non-zero. This question is studied in [Hu20] (in case $E=\mathbb{Q}$) for the abelian part by using some $\lambda$-independence results of $\mathbf{G}_{\lambda}$, such as the following.
@@ 3.2. Irreducibility and algebraic monodromy groups
Suppose $\det(\sigma_{\lambda})\cdot\tau_{\lambda}^{-2}$ is trivial. Then $\rho_{\lambda}\otimes\tau_{\lambda}^{-1}$ is self-dual. Since $\tau_{\lambda}$ is locally algebraic, $\pi$ is essentially self-dual by multiplicity one [JS81]. But in this case, $\rho_{\lambda}$ is irreducible (see [Hu23a, §$\mathsection 4.4.3$]). We conclude that $\rho_{\lambda}$ is irreducible.∎
Remark 3.1.
F. Calegari suggested to us an alternative proof of Theorem 1.2 that avoids the use of $\mathrm{Alt}^{2}(\rho_{\lambda})$. The key idea, in the case that $\rho_{\lambda}$ is reducible, was to add a suitable Hecke character $\chi$ to $\pi$, so that the isobaric sum $\pi\boxplus\chi$ is self-dual up to twist. His suggestion initially used Hodge-theoretic properties of $\rho_{\lambda}$ that currently do not seem to be in the literature unless $\pi$ is polarizable. However it can also be completed by using the results of the previous section as follows:
Suppose $\rho_{\lambda}$ is reducible with a decomposition $\rho_{\lambda}=\sigma_{\lambda}\oplus\tau_{\lambda}$ where $\sigma_{\lambda}$ is a two-dimensional representation and $\tau_{\lambda}$ is a character. As above, by Theorem 1.1, $\tau_{\lambda}$ comes from an algebraic Hecke character $\tau$ of $K$, and so $\tau_{\lambda}$ is $E^{\prime}$-rational for some number field $E^{\prime}\supset E$. Hence $\sigma_{\lambda}$ is $E^{\prime}$-rational, and this then also follows for $\det\sigma_{\lambda}$, so that again by Theorem 1.1, also $\det\sigma_{\lambda}$ comes from an algebraic Hecke character, say $\mu$, of $K$.
Define $\pi^{\prime}:=(\pi\otimes|\det|^{-1})\boxplus(\tau^{-1}\mu)$. Then one checks that $\rho_{\lambda}\oplus(\tau_{\lambda}^{-1}\det\sigma_{\lambda})$ has contragradient representation isomorphic to its twist by $(\det\sigma_{\lambda})^{-1}$, and it follows from multiplicity one that $(\pi^{\prime})^{\vee}\cong\pi^{\prime}\otimes\mu^{-1}$, i.e., that $\pi^{\prime}$ is essentially self-dual. Since $\pi^{\prime}$ is of isobaric type (3,1), $\pi$ is also essentially self-dual. We conclude by the last lines of the previous proof.∎
@@ Acknowledgments
G. Böckle and C.-Y. Hui would like to thank F. Calegari, T. Gee, C. Khare, M. Larsen, and J. Thorne for feedback and comments on this work. G. Böckle acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Collaborative Research Centre TRR 326 Geometry and Arithmetic of Uniformized Structures, project number 444845124. C.-Y. Hui was partially supported by NSFC (no. 12222120) and a Humboldt Research Fellowship.